# A Survey on Self-play Methods in Reinforcement Learning

### v1

@@ II-B3 Transitive Game and Non-transitive Game
For the sake of simplicity, we restrict our focus to two-player zero-sum symmetric games. In a transitive game, the strategies or outcomes adhere to a transitive relationship. Formally, $\forall\pi_{i},\pi_{j},\pi_{k}\in\Pi$, if $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, then it must follow that $u(\pi_{i},\pi_{k})>0$. This transitive property simplifies the strategic landscape, allowing for an ordinal ranking of strategies. Conversely, in a non-transitive game, $\exists\leavevmode\nobreak\ \pi_{i},\pi_{j},\pi_{k}\in\Pi$ such that $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, but $u(\pi_{i},\pi_{k})\leq 0$. This introduces a cyclic relationship among strategies, thereby complicating the game. The complexity often results in a mixed-strategy equilibrium, where players randomize their choices among multiple strategies to maximize their expected payoff. A quintessential example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others. In real-world settings, games exhibit complexities that extend beyond theoretical models. [29] argues that real-world games have two salient features: first, practice usually leads to performance improvements; and second, there are a plethora of qualitatively distinct strategies, each with unique advantages and disadvantages. In such games, the strategies form a geometric topology resembling a spinning top, where the vertical axis represents the performance of the strategy, and the radial axis represents the length of the longest cycle.
II-B4 Stage Game and Repeated Game
@@
A stage game (or one-shot game) is a game that is played only once, namely a one-shot interaction between players. A famous example of a stage game is the Prisoner’s Dilemma. A repeated game is derived from a stage game that is played multiple times. Formally, a repeated game based on a stage game $\mathcal{G}$ is defined by playing $\mathcal{G}$ for $T$ periods, where $T$ can be finite or infinite. The strategies in a repeated game are history-contingent, meaning that they can depend on the entire sequence of past plays. It’s important to note that a stage game or a repeated game can be either represented in the normal-form or extensive-form.
II-B5 Nash Equilibrium
@@
A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is a Nash equilibrium (NE) if, for every player $i$:
@@
meaning that no player can benefit by changing their strategy unilaterally, given the strategies of all other players. In other words, an NE is a situation where the strategy chosen by each player is a BR to the strategies chosen by all other players.
A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is an $\bm{\epsilon}$-NE if, for every player $i$:
@@
However, computing NE is generally intractable in complex games, leading some researchers to utilize $\alpha$-Rank [30] and Correlated Equilibrium (CE) [31] as alternatives. Additionally, some studies resort to Replicator Dynamics [32] as a method to analyze and understand the evolution of strategies within these games.
II-B6 Team Games
The framework of a two-player zero-sum game can be naturally extended to encompass team-based zero-sum games. Von Stengel and Koller analyzed zero-sum normal-form games involving a single team competing against an adversary [33]. In this type of team game, consider a team denoted by $\mathcal{T}=\{1,2,\cdots,n-1\}$. The player $n$ is the adversary $(\mathcal{D})$. In this kind of zero-sum normal-form team games, for any player $i,j\in\mathcal{T}$, the utility functions satisfy $u_{i}(\bm{\pi})=u_{j}(\bm{\pi})=u_{\mathcal{T}}(\bm{\pi})$ and $u_{\mathcal{D}}(\bm{\pi})=-(n-1)u_{\mathcal{T}}(\bm{\pi})$. A zero-sum single-team single-adversary normal-form game can also be extended to the domain of extensive games [34]. For any player $i,j\in\mathcal{T}$ and all terminal nodes $z\in Z$, the utility functions satisfy $u_{i}(z)=u_{j}(z)=u_{\mathcal{T}}(z)$ and $u_{\mathcal{D}}(z)=-(n-1)u_{\mathcal{T}}(z)$. Let $I_{\mathcal{T}}$ denote the information set defined as $\bigcup_{i\in\mathcal{T}}I_{i}$, and let $A_{\mathcal{T}}$ represent the set of actions accessible in the information sets within $I_{\mathcal{T}}$.
In scenarios where teammates are unable to coordinate their strategies, the team-maxmin equilibrium (TME) emerges as the most suitable solution concept [33]. We denote the collection of sequences of player $i$ by $Q_{i}$, representing the sequence-form actions undertaken by player $i$. The sequence-form strategy is encapsulated by a function $p_{i}:Q_{i}\rightarrow\mathbb{R}$, which maps each sequence $q\in Q_{i}$ to its associated probability of execution. Formally, the TME is articulated as:
$\arg\max_{p_{1},\ldots,p_{n-1}}\min_{p_{n}}u_{\mathcal{T}}\prod_{i=1}^{n}p_{i}.$ (5)
Similar to the arguments in normal-form games [33], it can also be inferred that in extensive-form games, a TME exists uniquely, barring any degeneracy and this TME aligns with the team’s utility maximization of the NE [34].
II-C Evaluation Metrics in Self-play
In this section, we introduce various self-play evaluation metrics, including NASHCONV (Section II-C1), Elo (Section II-C2), Glicko (Section II-C3), WHR (Section II-C4), and TrueSkill (Section II-C5). While NASHCONV measures the distance from Nash equilibrium, the other four metrics evaluate relative skill levels and are compared in Table I. It’s important to note that although numerous other evaluation metrics exist, the metrics highlighted here are among the most widely used in the field.
Table I: Comparison of Relative Skill Evaluation Metrics.
Elo Glicko WHR TrueSkill
Uncertainty Modeling $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Ratings At Any Time $\times$ $\times$ $\checkmark$ $\times$
Multiplayer In One Team $\times$ $\times$ $\times$ $\checkmark$
Bayesian Foundation $\times$ $\times$ $\checkmark$ $\checkmark$
II-C1 NASHCONV
Nash convergence (NASHCONV) serves as a metric to measure the deviation of a particular strategy from an NE. A lower NASHCONV value suggests that the strategy is closer to an NE, implying that no player would benefit from unilateral deviation from the strategy. Formally, it is defined as:
$\text{NASHCONV}(\bm{\pi})=\sum_{i}\max_{\pi_{i}\in\Pi_{i}}u_{i}(\pi_{i},\pi_{-i})-u_{i}(\bm{\pi}),$ (6)
where $\bm{\pi}$ denotes the combined strategy profile of all participating agents. In particular, in the context of two players, this deviation is commonly referred to as exploitability.
II-C2 Elo
The Elo system [35] operates under the assumption that the performance of each player in each game is a normally distributed random variable, with the mean being the player’s current rating. In a match between player A and player B, $R_{A}$ and $R_{B}$ are the current ratings of player A and player B. The probability density functions for the performance of player A and player B are given by:
$f_{A}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-R_{A})^{2}}{2\sigma^{2}}},$ (7)
$f_{B}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-R_{B})^{2}}{2\sigma^{2}}}.$ (8)
$E_{A}$ and $E_{B}$ denote the expected score (or the probability of winning) for player A and player B:
[…]
For convenience, the logistic function is used to approximate the probability (using the logistic curve with base 10 and choosing 400 to scale and normalize the difference in ratings):
$E_{A}$ $=\frac{1}{1+10^{(R_{B}-R_{A})/400}},$ (11)
$E_{B}$ $=\frac{1}{1+10^{(R_{A}-R_{B})/400}}.$ (12)
Note that $E_{A}+E_{B}=1$. $S_{A}$ is the actual outcome of the match for player A (1 for a win, 0.5 for a draw, 0 for a loss). $S_{B}=1-S_{A}$. After a match, the ratings are updated based on the difference between the actual outcome and the expected outcome. The adjustment is given by:
$\Delta R_{A}=K(S_{A}-E_{A}),$ (13)
$\Delta R_{B}=K(S_{B}-E_{B}).$ (14)
$K$ is a scaling factor, often determined by the specific domain of the application, and controls the maximum possible rating change for a single match. If two players have nearly identical ratings, the expected outcome will be close to 0.5. A win for either player will result in a moderate increase in their rating. In contrast, if there is a significant rating difference, the expected outcome will be heavily skewed. If the higher-rated player wins, their rating will increase by a value much smaller than $K$, reflecting the anticipated nature of their victory. However, if the lower-rated player secures an unexpected win, their rating will surge by a value approaching $K$, signifying the upset. The updated ratings become:
$R_{A}^{\prime}=R_{A}+\Delta R_{A},$ (15)
$R_{B}^{\prime}=R_{B}+\Delta R_{B}.$ (16)
However, there are challenges and limitations in Elo. First, the Elo system assumes that all matches are equally important. This might not be the case in all domains. Second, $K$ is often kept constant. However, a dynamic $K$ factor might be more appropriate, especially for players new to the rating pool or in scenarios where the importance of the match varies. Third, the standard Elo system does not incorporate a decay mechanism to account for the potential degradation or improvement of skills during periods of inactivity. Fourth, the basic Elo system is designed for one-on-one competitions. Adapting it to team or multi-player scenarios, such as team sports or online multiplayer games, can be challenging. Lastly, an important limitation of the Elo rating system is its unsuitability for games that exhibit high levels of non-transitivity [36].
II-C3 Glicko
The Glicko system refines the Elo system by introducing a measure of uncertainty or reliability in a player’s rating, termed rating deviation [37]. The primary motivation is to account for the variability in a player’s performance and the potential changes in skill over time. The Glicko-2 system, an extension of the original Glicko system, further refines these concepts and introduces the rating volatility $\sigma$, indicating the degree of expected fluctuation in the player’s rating [38].
In the Glicko-2 system, $r$ denotes the current rating of the player. $RD$ is the player’s rating deviation, and $\sigma$ is the player’s volatility. Convert the rating and rating deviation to the Glicko-2 scale:
$\mu$ $=\frac{r-1500}{173.7178},$ (17)
$\phi$ $=\frac{RD}{173.7178}.$ (18)
Then, calculate $v$ representing the estimated variance of the player’s rating based on game outcomes, $E(\mu,\mu_{j},\phi_{j})$ representing the probability of a player with rating $\mu$ defeating an opponent player $j$ and $\Delta$ representing the estimated improvement based only on game outcomes $s_{j}$:
$v$ $=\left[\sum_{j=1}^{m}g(\phi_{j})^{2}\{1-E(\mu,\mu_{j},\phi_{j})\}\right]^{-1},$ (19)
$E(\mu,\mu_{j},\phi_{j})$ $=\frac{1}{1+\exp(-g(\phi_{j})(\mu-\mu_{j}))},$ (20)
$\Delta$ $=v\sum_{j=1}^{m}g(\phi_{j})\{s_{j}-E(\mu,\mu_{j},\phi_{j})\},$ (21)
where:
$g(\phi)$ $=\frac{1}{\sqrt{1+3\phi^{2}/\pi^{2}}}.$ (22)
The update for $\sigma$ is more involved and requires an iterative procedure to solve for its new value $\sigma^{\prime}$. Then, calculate new $\mu^{\prime}$ and $\phi^{\prime}$:
$\mu^{\prime}$ $=\mu+\phi^{\prime 2}\sum_{j=1}^{m}g(\phi_{j})\{s_{j}-E(\mu,\mu_{j},\phi_{j})\},$ (23)
$\phi^{\prime}$ $=\frac{1}{\sqrt{\frac{1}{{\phi^{*}}^{2}}+\frac{1}{v}}},$ (24)
where:
$\phi^{*}=\sqrt{\phi^{2}+\sigma^{\prime 2}}.$ (25)
After calculations, the rating and rating deviation are converted back from the Glicko-2 scale:
$r^{\prime}$ $=173.7178\mu^{\prime}+1500,$ (26)
$RD^{\prime}$ $=173.7178\phi^{\prime}.$ (27)
II-C4 WHR
The Whole-History Rating (WHR) system [39] is a Bayesian rating system designed to estimate players’ skills from their entire game history. It is particularly adopted in handling the temporal dynamics of player skills. $R_{i}(t)$ denotes Elo rating of the player $i$ at time $t$. Similar to Equ. (11) and Equ. (12), $E_{A}(t)$ and $E_{B}(t)$ denote the expected score (or the probability of winning) for player A and player B at time $t$:
$E_{A}(t)$ $=\frac{1}{1+10^{(R_{B}(t)-R_{A}(t))/400}},$ (28)
$E_{B}(t)$ $=\frac{1}{1+10^{(R_{A}(t)-R_{B}(t))/400}}.$ (29)
Moreover, WHR assumes that ratings of each player $i$ is a Wiener process:
$r_{i}(t_{2})-r_{i}(t_{1})\sim\mathcal{N}(0,|t_{2}-t_{1}|\omega^{2}),$ (30)
where $\omega$ is a parameter representing the variability of ratings in time. Thus, if $r^{\prime}=R_{A}(t_{1}),r^{\prime\prime}=R_{A}(t_{2})$:
$p(r^{\prime\prime}|r^{\prime})=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2}(\frac{r^{\prime\prime}-r^{\prime}}{\sigma})^{2}},$ (31)
where $\sigma=\omega\sqrt{|t_{2}-t_{1}|}$. In addition, a Wiener process is memoryless (or Markovian).
$\mathbf{R}(t)$ denotes player A’s ratings, $\mathbf{G}$ denotes the observation of game results. Thanks to the Bayes formula, the WHR system is to solve:
$\arg\max_{\mathbf{R}}p(\mathbf{R}|\mathbf{G})=\arg\max_{\mathbf{R}}\frac{p(\mathbf{G}|\mathbf{R})p(\mathbf{R})}{p(\mathbf{G})}.$ (32)
$P(\mathbf{G}|\mathbf{R})$ can be derived from the product of Equ. (28) and $p(\mathbf{R})$ can be derived from the product of Equ. (31). Utilizing the iterative process of Newton’s method, we can ascertain the solution to the given problem.
II-C5 TrueSkill
TrueSkill [40] is based on a probabilistic graphical model that employs Bayesian inference and adapts to multiple players in multiple teams. TrueSkill 2 [41] extends the original TrueSkill model with several enhancements by taking into account the experience of a player, the affiliation with a team, and some game-specific factors such as the number of kills. For simplicity, TrueSkill is introduced using the following scenario: team 1 has player 1, while team 2 has player 2 and player 3. Ultimately, team 1 defeated team 2. This is a simpler case than the one presented in the original paper [40]. It can be described in an undirected probabilistic graphic model (Fig. 2). The skill of each player, denoted by $s_{i}$, is represented by a Gaussian distribution $N(s_{i};\mu_{i},var_{i})$. The performance of each player, denoted by $p_{i}$, is also represented by a Gaussian distribution $N(p_{i};s_{i},\beta^{2})$. The performance of each team is denoted by $t_{i}$. The difference in team performance is denoted by $d$. If $d>\epsilon$, where $\epsilon$ is a small positive threshold, then team 1 beats team 2. The TrueSkill algorithm’s updating mechanism within the context of undirected probabilistic graphs uses the sum-product message-passing algorithm. This ensures that the skill estimates are refined iteratively, leading to accurate and reliable ratings for each player.
Figure 2: An undirected probabilistic graphic model example of TrueSkill.
III Algorithms
Based on existing self-play work [18, 42, 43, 44], we propose a self-play framework (Algo. 1) that boasts enhanced expressivity and superior generalization capabilities. The framework is adept at handling multi-homogeneous-player general-sum games. It’s important to note that although homogeneous players represent a specific subset of heterogeneous players, the latter can be reformulated as the former by expanding the dimensions of the input vector, which essentially entails embedding agent identity information. Moreover, as players are homogeneous, we assume each player shares the same policy population.
To clarify our framework, we describe it in a more easily understandable way. All players share a common policy population with a fixed maximum size. In each iteration, a new policy is initialized for training, and opponent policies are sampled from the existing policy population. During this iteration, the opponent policies typically remain fixed, while only the policy being trained is updated. After the training process, the new policy replaces one of the policies in the policy population. An evaluation metric is then used to assess the performance of the updated policy population. Based on this performance, the strategy for sampling opponents is adjusted for the next iteration. This process is repeated. For a more detailed and precise description, please refer to Sec. III-A.
In addition, we divide self-play algorithms into four primary groups: traditional self-play algorithms (Sec. III-B), the PSRO series (Sec. III-C), the ongoing-training-based series (Sec. III-D), and the regret-minimization-based series (Sec. III-E). We analyze how these four categories align with our framework in their respective sections and introduce representative algorithms in each category. Moreover, in Sec. III-F, we discuss the differences among these four categories and their connections to our proposed framework. We also explain why our framework is more general than existing frameworks. Furthermore, we summarize the major elements of key algorithms within our framework for each category in Table II.
III-A Framework Definition
Algorithm 1 the Framework of Self-play
0:  $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$$\triangleright$ Policy population.
0:  $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$$\triangleright$ Interaction matrix.
0:  $\mathcal{F}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$$\triangleright$ MSS.
1:  for $e\in[[E]]$ do
2:     for $\sigma_{i}\in\Sigma$ do
3:        Initialize $\pi_{i}^{h}$
4:        $\pi_{i}^{h}\leftarrow\text{ORACLE}(\pi_{i}^{h},\sigma_{i},\Pi)$$\triangleright$ Compute the oracle.
5:        $\mathcal{P}\leftarrow\text{EVAL}(\Pi)$triangleright Get policies’ performance.
6:        $\Sigma\leftarrow\mathcal{F}(\mathcal{P})$$\triangleright$ Update the interaction matrix.
7:     end for
@@ III-A Framework Definition
9:  return  $\Pi,\Sigma$
In Algo. 1, we define a unified framework of self-play based on [43, 42, 44, 18]. First, the input of the framework is defined as follows:
  • $\Pi$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which is determined by specific algorithms. We will further introduce this policy condition function in the following sections. For simplicity, we denote each policy $\pi_{i}(\cdot|h(i))$ by $\pi_{i}^{h}$. It’s important to note that $i$ refers to the $i$th policy in the policy population, not to a player. Additionally, the hyper-parameter $N$ denotes the policy population size of $\Pi$, not the number of game players. Furthermore, $\Pi$ can be initialized in two different ways. First, $\Pi$ can be considered as initialized with $N$ placeholder policies, meaning that there is no actual initialization of the policies in practice. Instead, each policy will be actually initialized in the training iterations (Line 3 in Algo. 1). We refer to this approach as the placeholder initialization. Second, $\Pi$ can be initialized with $N$ real policies, which may include random initialization or pretrained models. We refer to this approach as the actual initialization. A policy $\pi_{i}^{h}$ is considered an ineffective policy if it serves as a placeholder policy; otherwise, it is deemed an effective policy. Additionally, Table II provides a clearer illustration of the initialization of $\Pi$ and the expression of $h(i)$ across different categories of the main self-play algorithms.
  • $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$: The interaction matrix of the policy population. $\sigma_{i}\in\mathbb{R}^{C_{1}}$ represents the opponent sampling strategy of policy $i$. Namely, $\sigma_{i}$ illustrates how to sample opponent(s)’ policies against policy $i$. For instance, let $\sigma_{i}$ represent the probability of policies for each opponent. In this context, $C_{1}=N^{n-1}$, where n denotes the number of players. Alternatively, $\sigma_{i}$ can also be considered as sampling parameters within a sampling network. Especially, in a two-player game, if $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, $\Sigma$ can be depicted in directed interaction graphs (Fig. 3). It’s important to note that, unlike the original PSRO framework [42], $\sigma_{i}$ here is not the meta-strategy of policy $i$. Instead, in the two-player setting, if $\sigma_{i}$ in our framework is the opponent’s meta-strategy against policy i, our framework can reduce to the original PSRO framework.
  • $\mathcal{F}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$: A meta-strategy solver (MSS) $\mathcal{F}$ takes the performance matrix $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$ as its input and produces a new interaction matrix $\Sigma\in\mathbb{R}^{N\times C_{1}}$ as its output. $p_{i}$ is the performance of policy i. For instance, $p_{i}$ can be depicted as the relative skill like Elo ratings ($C_{2}=1$) or can be depicted as the payoff tensor ($C_{2}=N^{n-1}$, where n is the number of players). Specially, in two-player symmetric zero-sum games, the expected payoffs can serve as the evaluation metric. In such cases, $\Sigma$ is a square matrix ($C_{2}=N$). In addition, Table II concludes the MSS of representative self-play algorithms across these four categories.
@@
Figure 3: When $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, we can consider various examples of conventional self-play algorithms: In the Top section, we define $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$. In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one. The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $i$ to node $j$ with a weight of $\sigma_{ij}$ signifies that policy $i$ is optimized against policy $j$ with a probability of $\sigma_{ij}$. Note that this figure is reproduced from [43] and this concept is initially proposed by [44].
Next, the core process within the framework is as follows:
@@
  • For epoch $e\in[[E]]$ (Line 1 in Algo. 1): $E$ denotes the total number of epochs for the entire policy population. For example, if the algorithm only introduces new policies into the population without updating existing ones, then $E=1$. This implies that only the ineffective policy is likely to be chosen for training in each iteration and turned into an effective one, while effective policies remain unchanged. Conversely, if effective policies are updated multiple times throughout the algorithm, then $E>1$. Indeed, $E$ accurately rechoflects the number of updates performed. Additionally, Table II summarizes the values of $E$ for different categories of self-play algorithms.
@@
  • Initialize $\pi_{i}^{h}$ (Line 3 in Algo. 1): The initialization of $\pi_{i}^{h}$ can vary depending on the algorithm being used. For instance, it may be initialized randomly, by leveraging pre-trained models [8, 15] or through a recently updated policy. We provide detailed descriptions of the initialization process for each algorithm series. Also, these are summarized in Table II.
@@
  • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 4 in Algo. 1): The ORACLE is an abstract computational entity that returns a new policy adhering to specific criteria. Here, we divide ORACLE into three types. (1) One type is the BR oracle, which is designed to identify the optimal counter-strategies against an opponent’s strategy, including finding NE [45]. However, it often requires considerable computational effort. This computational demand can be a limitation, particularly in complex environments or with large action spaces. (2) To alleviate the computational demands, the approximate best response (ABR) oracle is introduced, which can be calculated using techniques such as RL (Algo. 2), evolution-theory-based methods (Algo. 3) or regret minimization methods (Algo. 4). (3) Additionally, other specially crafted ORACLEs are either tailored to new MSSes [46, 47] or introduced to enhance diversity [48, 49, 50].

### v2

@@ 2.2.3 Transitive Game and Non-transitive Game
For the sake of simplicity, we restrict our focus to two-player zero-sum symmetric games. In a transitive game, the strategies or outcomes follow a transitive relationship. Formally, $\forall\pi_{i},\pi_{j},\pi_{k}\in\Pi$, if $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, then it must follow that $u(\pi_{i},\pi_{k})>0$. This transitive property simplifies the strategic landscape, allowing for an ordinal ranking of strategies. Conversely, in a non-transitive game, $\exists\leavevmode\nobreak\ \pi_{i},\pi_{j},\pi_{k}\in\Pi$ such that $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, but $u(\pi_{i},\pi_{k})\leq 0$. This introduces a cyclic relationship among strategies, thereby complicating the game. The complexity often results in a mixed-strategy equilibrium, where players randomize their choices among multiple strategies to maximize their expected payoff. A classical example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others.
2.2.4 Stage Game and Repeated Game
@@
A stage game (or one-shot game) is a game that is played only once, namely a one-shot interaction between players. A famous example of a stage game is the Prisoner’s Dilemma. A repeated game is derived from a stage game played multiple times. Formally, a repeated game based on a stage game $\mathcal{G}$ is defined by playing $\mathcal{G}$ for $T$ periods, where $T$ can be finite or infinite. The strategies in a repeated game are history-contingent, meaning they can depend on the entire sequence of past plays. An example of a repeated game is Texas Hold’em, where players engage in multiple rounds, and each player’s strategy may evolve based on previous rounds, betting patterns, and the history of interactions among players.
2.2.5 Nash Equilibrium
@@
A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is a Nash equilibrium (NE) if, for each player $i$:
@@
meaning that no player can benefit by changing their strategy unilaterally, given the strategies of others. In other words, an NE is a situation where each player’s strategy is a BR to the others’ strategies.
A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is an $\bm{\epsilon}$-NE if, for each player $i$:
@@
However, computing NE is generally intractable in complex games, so some researchers utilize $\alpha$-Rank [101] and Correlated Equilibrium (CE) [2] as alternatives. Some studies also resort to Replicator Dynamics [124] to analyze strategy evolution.
3 Algorithms
Based on existing self-play work [54, 73, 82, 42], we propose a self-play framework (Algo. 1) that boasts enhanced expressivity and superior generalization capabilities. For simplicity, our framework is illustrated for symmetric games. All players share a policy population with a fixed maximum size. In each iteration, a newly initialized policy is trained, while opponent policies are sampled from the population. After training, the new policy is added to the population. The updated policy population is then evaluated, and then the opponent sampling strategy is recalculated for the next iteration.
This section is organized as follows: In Sec.3.1, we provide a formalized description of our framework. We then categorize self-play algorithms into four primary groups: traditional self-play algorithms (Sec. 3.2), the PSRO series (Sec. 3.3), the ongoing-training-based series (Sec. 3.4), and the regret-minimization-based series (Sec. 3.5). We analyze how these four categories align with our framework and introduce corresponding algorithms for each category. To make it more straightforward, we highlight the representative classic algorithms from these and present them in Table 1 for comparison within our framework. In Sec. 3.6, we compare the four categories, explain our classification rationale, and highlight the distinctions between our framework and existing ones to demonstrate its greater generality.
3.1 Framework Definition
Figure 3: Illustration of our framework. The top row shows two different initialization methods for the policy population $\Pi$ and the interaction matrix $\Sigma$: lazy initialization and immediate initialization (corresponding to Line 1 in Algo. 1). The bottom row illustrates two training paradigms: standard training and ongoing training (corresponding to Lines 2~9 in Algo. 1). We categorize self-play algorithms into four types: traditional self-play, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. Among these, traditional self-play, the PSRO series, and the regret-minimization-based series utilize lazy initialization and standard training, whereas the ongoing-training-based series employs immediate initialization and ongoing training. For a more detailed description and analysis, please refer to Sec.3 and Table1.
Algorithm 1 the Framework of Self-play.
1:  Initialize $\Pi,\Sigma$
2:  for $e\in[[E]]$ do
3:     for $\sigma_{i}\in\Sigma$ do
4:        Initialize $\pi_{i}^{h}$
5:        $\pi_{i}^{h}\leftarrow\textit{ORACLE}(\pi_{i}^{h},\sigma_{i},\Pi)$$\triangleright$ Compute the oracle.
6:        $\mathcal{P}\leftarrow\textit{EVAL}(\Pi)$triangleright Get policies’ performance.
7:        $\Sigma\leftarrow\textit{MSS}(\mathcal{P})$$\triangleright$ Update the interaction matrix.
@@ 3.1 Framework Definition
9:  end for
10:  return  $\Pi,\Sigma$
In Algo. 1, we define a unified self-play framework based on [82, 73, 42, 54]. We combine visual illustrations (Fig. 3) with detailed textual descriptions to explain the key processes of our framework. Next, we provide an in-depth explanation of these processes:
@@
Figure 4: Interaction matrix $\Sigma$ examples. When $C_{1}=N$ and the opponent sampling strategy $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, the interaction matrix $\Sigma$ can be depicted in directed interaction graphs. Here, we consider three representative self-play algorithms. In the Top section, we define the interaction matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$. In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one. The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $i$ to node $j$ with a weight of $\sigma_{ij}$ signifies that policy $i$ is optimized against policy $j$ with a probability of $\sigma_{ij}$. Note that this figure is reproduced from [82] and this concept is initially proposed by [42].
  • $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which provides supplementary information for certain algorithms. We denote the policy $\pi_{i}(\cdot|h(i))$ as $\pi_{i}^{h}$. Note that $i$ refers to the $i$th policy in the population, not the $i$th player, and $N$ denotes the policy population size, not the number of players.
    $\Pi$ can be initialized (Line 1 in Algo. 1) in two ways: lazy initialization and immediate initialization. In lazy initialization, $\Pi$ starts with $N$ placeholder policies, with the actual policies being initialized during the training iterations (Line 4 in Algo. 1). In immediate initialization, $\Pi$ is initialized upfront with $N$ real policies, which may be randomly generated or pre-trained models. A policy $\pi_{i}^{h}$ is considered inactive if it is a placeholder; otherwise, it is considered active.
  • $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$: The interaction matrix $\Sigma$ consists of rows $\sigma_{i}$, where each $\sigma_{i}\in\mathbb{R}^{C_{1}}$ represents the opponent sampling strategy of policy $i$, indicating how to sample opponents’ policies. The dimension $C_{1}$ varies according to the sampling method. For instance, $\sigma_{i}$ can represent the probability distribution over opponent policies in the policy population $\Pi$, where $C_{1}=N^{n-1}$ and $n$ denotes the number of players. Alternatively, $\sigma_{i}$ can be viewed as parameters of a sampling network. Specially, in a two-player game, if $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, $\Sigma$ can be depicted in directed interaction graphs (Fig. 4).
    $\Sigma$ will be initialized (Line 1 in Algo. 1) based on the initialization of $\Pi$. If $\Pi$ uses lazy initialization, $\Sigma$ will be initialized as an empty matrix. If $\Pi$ uses immediate initialization, $\Sigma$ will be calculated based on $\Pi$. We will discuss this in more detail later.
@@
  • For epoch $e\in[[E]]$ (Line 2 in Algo. 1): $E$ denotes the total number of epochs for the entire policy population. For example, if the algorithm only introduces new policies into the population without updating existing ones, then $E=1$. This implies that only the inactive policy is likely to be chosen for training in each iteration and turned into an active one, while activa policies remain unchanged. Conversely, if active policies are updated multiple times throughout the algorithm, then $E>1$. Indeed, $E$ accurately reflects the number of updates performed.
@@
  • Initialize $\pi_{i}^{h}$ (Line 4 in Algo. 1): The initialization of $\pi_{i}^{h}$ can vary depending on the algorithm being used. It can be initialized randomly or by leveraging pre-trained models [115, 129] or through a recently updated policy.
@@
  • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 5 in Algo. 1): The ORACLE is an abstract computational entity that returns a new policy adhering to specific criteria. Here, we divide ORACLE into three types. (1) One type is the BR oracle, which is designed to identify the optimal counter-strategies against an opponent’s strategy, including finding NE [92]. However, it often requires considerable computational effort. (2) To alleviate the computational demands, the approximate best response (ABR) oracle is introduced, using techniques such as RL (Algo. 2), evolution-theory-based methods (Algo. 3) or regret minimization methods (Algo. 4). (3) Other specially crafted ORACLEs are either tailored to specific meta-strategy solvers (MSSes) [96, 86] or introduced to enhance diversity [4, 103, 83].